# What’s Wrong with Your Code Generated by Large Language Models? An Extensive Study

### v1

@@
Abstract.
The increasing development of large language models (LLMs) in code generation has drawn significant attention among researchers. To enhance LLM-based code generation ability, current efforts are predominantly directed towards collecting high-quality datasets and leveraging diverse training technologies. However, there is a notable lack of comprehensive studies examining the limitations and boundaries of these existing methods. To bridge this gap, we conducted an extensive empirical study evaluating the performance of three leading closed-source LLMs and four popular open-source LLMs on three commonly used benchmarks. Our investigation, which evaluated the length, cyclomatic complexity and API number of the generated code, revealed that these LLMs face challenges in generating successful code for more complex problems, and tend to produce code that is shorter yet more complicated as compared to canonical solutions. Additionally, we developed a taxonomy of bugs for incorrect codes that includes three categories and 12 sub-categories, and analyze the root cause for common bug types. Furthermore, to better understand the performance of LLMs in real-world projects, we manually created a real-world benchmark comprising 140 code generation tasks. Our analysis highlights distinct differences in bug distributions between actual scenarios and existing benchmarks. Finally, we propose a novel training-free iterative method that introduces self-critique, enabling LLMs to critique and correct their generated code based on bug types and compiler feedback. Experimental results demonstrate that our approach can significantly mitigate bugs and increase the passing rate by 29.2% after two iterations, indicating substantial potential for LLMs to handle more complex problems.
@@ Abstract.
Shihan Dou and Haoxiang Jia contributed equally to this paper. Ming Wen and Tao Gui are the corresponding authors.
1. Introduction
@@
Code generation (or program synthesis) aims to accurately generate executable programs from problem specifications such as natural language requirement descriptions and unit test examples(Le et al., 2022; Svyatkovskiy et al., 2020). Research in code generation has gained considerable attention due to its significant impact on the software industry, particularly in enhancing productivity and accessibility by accelerating program development and improving code reliability (Shin and Nam, 2021; Dehaerne et al., 2022; Zan et al., 2022).
@@
Recently, the advancement of large language models (LLMs) has significantly propelled the field of code generation (Wang et al., 2021; Feng et al., 2020; Zhang et al., 2023; Lu et al., 2021). Trained on large corpora of text and code, LLMs can understand human requirements and generate corresponding code through “next code token prediction” (Radford et al., 2019, 2018; Brown et al., 2020). However, the existing LLMs struggle to generate correct code in the face of more complicated human requirements (Dou et al., 2024; Roziere et al., 2023; Hendrycks et al., 2021; Li et al., 2023). Despite efforts to enhance code generation accuracy using diverse pre-trained code corpora (Guo et al., 2024; Lozhkov et al., 2024; Song et al., 2024), fine-tuning models on higher-quality code (Zheng et al., 2024; Wei et al., 2024), or employing various natural language process (NLP) techniques (Wang et al., 2022c; Yu et al., 2023; Zhou et al., 2022; Wang et al., 2023), there remains a lack of a thorough empirical study on “what’s wrong with the code generated by large language models”.
First, accuracy (i.e., passing rate of unit tests) is by far the most dominant code generation criterion, encompassing both syntactic and semantic requirements (Yuan et al., 2023; Fix, 2009). Other characteristics such as code length and complexity are equally significant due to the importance of code maintenance (Alawad et al., 2019; Tashtoush et al., 2023). Therefore, we expect the code generated by LLMs to be precise, with short code length and low code complexity. Second, annotating the bug types in generated code and analyzing their distribution is necessary. There are a multitude of reasons why code generated by LLMs may not reach human requirements. During code generation, LLMs must fully comprehend the problem description and generate the corresponding code (Jain et al., 2023; Liu et al., 2023b; Wong et al., 2023; Murr et al., 2023). However, deviations in the LLM’s understanding of the actual intent can lead to task failure. Additionally, the generated code might contain syntactic or semantic errors. The generation paradigm of LLMs, which is the pattern of predicting the next code token, could also introduce formatting errors as it is driven by probability and does not inherently ensure consistency in indentation and syntax (Le et al., 2022; Dou et al., 2024). By categorizing bugs and analyzing their distributions, we can comprehensively understand the limitations of LLMs in code generation tasks. This analysis is crucial for developing strategies to address these bugs, thereby improving the quality and accuracy of generated code. Finally, there is a noticeable deficiency in benchmarks for evaluating LLMs in real-world scenarios. The benchmarks commonly used are constructed from programming questions (Austin et al., 2021; Hendrycks et al., 2021; Chen et al., 2021) or derived from real-world code repositories and libraries (Zan et al., [n. d.]; Li et al., 2024). These same repositories and libraries are used for pre-training LLMs, which could cause data contamination and test-case leakage (Jain et al., 2024; Golchin and Surdeanu, 2023; Weller et al., 2023; Chen et al., 2021; Riddell et al., 2024; Roberts et al., 2023). Therefore, developing a high-quality, previously untrained benchmark to evaluate LLMs under real-world conditions is crucial.
@@
In this paper, we conduct a comprehensive large-scale empirical study to explore the capabilities and limitations of code generation using large language models. We thoroughly examine the correctness and complexity of the generated code. We categorize bugs using script-based initial annotation and detailed manual annotation, statistically modeling their distribution and providing an in-depth analysis. To assess the effectiveness of LLMs in real-world code generation, we manually constructed a high-quality benchmark, RWPB, and evaluated LLM performance on it. Based on our findings, we propose a novel, practical method introducing self-critique to iteratively identify and fix bugs in LLM-generated code without additional training.
Our study aims to answer the following research questions:
@@
  • RQ2: Bugs in Code Generated by LLMs (§4). What are the different types of root causes (i.e., bugs) that cause incorrect code to be generated? How are these bugs distributed across widely used LLMs and popular benchmarks?
@@
  • RQ3: Benchmark Construction and Evaluation (§5). How to build a real-world benchmark while minimizing data leakage (i.e., ensuring it is not used in pre-training). How do real-world project bugs differ from common, widely used benchmarks?
@@
  • RQ4: Mitigating Bugs in Generated Code (§6). How to fix bugs based on our findings without introducing additional resource consumption, such as fine-tuning LLMs.
@@
To investigate these RQs, we collected 1,164 programming problems from three widely used benchmarks (HumanEval+ (Chen et al., 2021), MBPP+ (Austin et al., 2021), and APPS+ (Hendrycks et al., 2021)). To perform code generation, we selected seven LLMs, including three state-of-the-art closed-source LLMs (GPT-4 (gpt, 2023b), GPT-3.5 (gpt, 2023a), and Claude-3 (Anthropic, 2024)), two well-known open-source LLMs (LLama-3-Instruct (lla, 2024) and Phi-3-Instruct (Abdin et al., 2024)), and two popular open-source code LLMs (StarCoder-2 (Lozhkov et al., 2024), DeepSeekCoder (Guo et al., 2024)). We evaluated these LLMs on generated code by using unit tests. Experimental results show that these LLMs could achieve on average a passing rate of 41.6%. Specifically, GPT-4 and Claude-3 demonstrate the best performance, which is 63.8% and 56.7%, respectively, while Phi-3 has the lowest accuracy at 30.9%.
@@
Furthermore, we collected the token number of the problem description, the length, the cyclomatic complexity and the API number of the canonical solution. This investigation allows us to analyze how the length of problem descriptions and the complexity of canonical solutions affect each model’s performance. We also examined the characteristics of the generated code, focusing on factors such as code characteristics and comment numbers across different LLMs and benchmarks.
@@
Additionally, to explore the root causes of incorrect code generated by LLMs, we conducted a comprehensive study on the bug taxonomy in generated code. Specifically, this analysis included two key steps: (1) First, we utilized scripts to automatically create an initial taxonomy of potential bugs based on compiler feedback. (2) Second, human experts manually examined the incorrect code and annotated the bug type. During this phase, experts were given the flexibility to expand and refine the bug taxonomy as needed. This process involved 22 human experts with backgrounds in software engineering, with at least two experts assigned to double-check each annotation. Ultimately, we classify the bugs into three primary bug types and 12 secondary bug types. Based on the above result, some notable findings are: (i) Closed-source models outperform open-source models, particularly when handling more complex tasks. (ii) Functional bugs constitute the highest proportion, whereas syntax bugs represent the lowest. (iii) The code generated by the models typically has fewer lines but higher complexity and a similar number of APIs as compared to the canonical solutions. (iv) Although generating more comments during code generation can enhance LLMs’ reasoning abilities, LLMs still struggle to solve more complicated problems. (v) For complex problems, LLMs often fail to generate optimal algorithms, leading to timeout errors.
To evaluate the efficacy of LLM code generation in real-world projects, we developed a real-world benchmark by manually collecting 140 code generation tasks from GitHub repositories established in 2024. Our experimental results revealed that closed-source LLMs outperform open-source LLMs in code generation quality for real-world projects, particularly excelling in reducing syntax bugs and runtime bugs. Claude-3 achieves the best performance with an accuracy of 45.7%, while Phi-3 has the lowest accuracy at only 22%. We also found that the distributions of bug taxonomy are inconsistent between the existing commonly used benchmarks and the real-world projects.
@@
Our study revealed that existing LLMs are prone to introducing bugs in code generation due to their failure to capture and interpret human intent and perform the correct corresponding logic, evidenced in both current popular benchmarks and real-world projects. To mitigate these bugs, we propose an interactive approach involving self-critique (Tan et al., 2023) without additional training costs. Specifically, we make LLMs critique their incorrect code based on our constructed bug taxonomy and compiler feedback (i.e., error messages from failed executions) to understand the problem in-depth and analyze the causes of the bug, and then re-correct the code. This method can be performed iteratively. Experimental results showed that our proposed method can mitigate bugs and increase the passing rate by 29.2% after two iterations, which indicates substantial potential for LLMs to handle more complex tasks.
@@
  • A thorough evaluation of the capabilities of LLMs in code generation. We conducted extensive experiments to evaluate the performance of current widely used LLMs. We also explored the effect of different types of tasks on code generation and the characteristics of generated code.
  • A bug taxonomy of LLM-generated code. We developed a bug taxonomy based on code generated by LLMs. We further analyzed these bugs and modeled their distribution to find the root causes of inaccuracies in LLM-generated code.
  • A real-world benchmark. We manually constructed a real-world benchmark by collecting code on the newest GitHub repositories. We also analyzed and compared the distribution of bugs between existing benchmarks and our real-world benchmark, highlighting the inadequacies of current models in real-world scenarios.
  • A novel method to mitigate code generation bugs. We propose a novel approach to bug mitigation by introducing self-critique: allowing LLMs to find and fix bugs by iteratively critiquing their generated code. We also offer some suggestions and possible future solutions for improvement111 Our labelling results, evaluating scripts, and real-world benchmark are available in our artifact repository (ar2, 2023)..
Table 1. Overview of seven LLMs used in our study. C.W. denotes Context Window.
Modality Model C.W. Release Date
@@
Table 2. Modalities of three common-used benchmarks and performance of seven popular LLMs in code generation on these benchmarks. $C$ denotes Cyclomatic Complexity. All values are in %.
Dataset Token Nums of Description Lines of Code CC of Codes API Nums of Code Open-Source Closed-Source Average
StarCoder-2 DeepSeekCoder Llama-3 Phi-3 GPT-4 GPT-3.5 Claude-3
HumanEval Plus 131.3 11.3 3.1 3.4 61.6 58.5 51.8 67.1 85.4 64.6 79.3 64.8
MBPP Plus 48.7 6.1 2.2 1.8 63.2 66.4 61.1 63.2 79.2 73.9 74.9 67.9
APPS Plus Introductory 307.7 10.3 3.2 5.6 42.0 45.5 37.5 13.5 72.0 56.0 69.5 44.4
Interview 480.3 27.4 6.4 15.1 10.5 20.5 17.5 9.0 46.5 25.0 38.0 21.5
Competition 554.9 33.7 6.9 18.4 2.0 1.0 7.5 1.5 36.0 8.5 22.0 9.4
Average 447.6 23.8 5.5 13.0 18.2 22.3 20.8 8.0 51.5 29.8 43.2 25.1
Average 304.6 17.8 4.4 8.9 35.9 38.4 35.1 30.9 63.8 45.6 56.7 41.6
2. Experimental Design
Large Language Model Selection. To thoroughly evaluate LLM-based code generation, we selected seven large language models of three types: open-source text LLMs, open-source code LLMs, and closed-source LLMs. For open-source text LLMs, we selected two common models, namely Llama-3 (lla, 2024) and Phi-3 (Abdin et al., 2024). Llama-3 is an advanced and popular LLM developed by Meta-AI, which is trained on extensive text and code corpora. Phi-3 is another extensively used LLM, noted for its training on high-quality textbook datasets. In the category of open-source code LLMs, our study used StarCoder-2 (Lozhkov et al., 2024) and DeepseekCoder (Lozhkov et al., 2024). StarCoder-2 is a widely used code model trained on 3.3 to 4.3 trillion tokens of code from The Stack v2 dataset (Lozhkov et al., 2024), encompassing over 600 programming languages. DeepseekCoder is another state-of-the-art open-source code model. Its training corpus is comprised of 2 trillion tokens, combining code and natural languages. We also selected three universally used closed-source LLMs, namely GPT-3.5 (gpt, 2023a), GPT-4 (gpt, 2023b), and Claude-3 (Anthropic, 2024). Additional information and details on these models are provided in Table 1.
@@ 2. Experimental Design
Benchmark Selection. To evaluate the accuracy and behavior of the generated code and to analyze their bugs, we constructed a collection of assessments from well-known benchmarks. Our selection criteria were as follows: benchmarks must be widely recognized to ensure relevance; they should not originate from popular libraries or projects to prevent data leakage, as current LLMs are trained on a significant portion of such data (Jain et al., 2024; Chen et al., 2021; Oren et al., 2023; Weller et al., 2023; Riddell et al., 2024; Roberts et al., 2023); and they must have diverse unit tests with high code coverage for the sample. Moreover, we selected benchmarks requiring the target programming language of the generated code to be Python, as most current research aiming to improve a model’s code-generation ability utilizes Python (Guo et al., 2024; Contributors, 2023; Zhuo et al., 2024). Finally, we selected three widely used benchmarks, HumanEval+ (Chen et al., 2021), MBPP+ (Austin et al., 2021) and APPS+ (Hendrycks et al., 2021), and constructed our collection of assessments based on them. HumanEval is a popular benchmark comprising 164 hand-written Python interview problems designed to test language comprehension, algorithmic thinking, and basic mathematics skills. MBPP is another prevalent benchmark containing 974 introductory instances created by crowd-sourced developers knowledgeable in Python. Both HumanEval+ and MBPP+ are expansions of their original versions, augmenting additional unit tests (Liu et al., 2024b). APPS is a code generation benchmark collected from open-access sites, including Codewars, AtCoder, Kattis, and Codeforces. APPS+ is a curated version of APPS that excludes instances lacking input, output, or canonical solutions, and includes manual analysis to eliminate issues such as incomplete code, syntax errors, API misuse, and missing library dependencies (Dou et al., 2024). Our benchmark includes all samples from HumanEval+ and MBPP+, along with 600 samples from APPS+, totaling 1,000 samples. Specifically, APPS+ categorizes samples into three levels of difficulty: introductory, interview, and competition. We randomly selected 200 samples from each category, with a total of 600 samples, to ensure a balance of varying levels of difficulty.
@@
Additionally, to evaluate the efficacy of LLM code generation in real-world projects, we created a real-world benchmark RWPB from GitHub repositories established in 2024 (details in Section 5.1).
@@
Prompt Design. To ensure experimental consistency across various LLMs, we employed a standardized prompt format, as outlined in (Liu et al., 2024b). Notably, our prompts excluded library import information, necessitating that the models autonomously identify and import the required libraries for the APIs utilized.
Task Prompt: Please provide a self-contained Python script that solves the following problem in a markdown code block: {Programming problem} Below is a Python script with a self-contained function that solves the problem and passes corresponding tests:
@@
Computation and Labor Resources. To evaluate current large language models (LLMs) for code generation and validate our proposed improvements, we conducted experiments on a single node with eight A100-80GB GPUs and 2TB of CPU memory, using Python 3.11. For code evaluation, bug analysis, and the construction of our real-world project benchmark (RWPB), we assigned 37 software engineering annotators, with backgrounds in software engineering. Given their countries of residence, all annotators obtained adequate compensation.
3. LLM-Based Code Generation

### v2

# Dou S H. What is Wrong with Your Code Generated by Large Language Models? An Extensive Study

@@
Abstract
The increasing development of large language models (LLMs) in code generation has drawn significant attention among researchers. To enhance LLM-based code generation ability, current efforts are predominantly directed towards collecting high-quality datasets and leveraging diverse training technologies. However, there is a notable lack of comprehensive studies examining the limitations and boundaries of existing methods. To bridge this gap, we conducted an extensive empirical study evaluating the performance of three leading closed-source LLMs and six popular open-source LLMs on three commonly used benchmarks. Our investigation, which evaluated the length, cyclomatic complexity and API number of the generated code, revealed that these LLMs face challenges in generating successful code for more complex problems, and tend to produce code that is shorter yet more complicated as compared to canonical solutions. Additionally, we developed a taxonomy of bugs for incorrect codes that includes three categories and ten sub-categories, and analyzed the root cause for common bug types. To better understand the performance of LLMs in real-world projects, we also manually created a real-world benchmark RWPB. We analyzed bugs on RWPB to highlight distinct differences in bug distributions between actual scenarios and existing benchmarks. Finally, we propose a novel training-free iterative method that introduces self-critique, enabling LLMs to critique and correct their generated code based on bug types and compiler feedback. Experimental results demonstrate that our approach can significantly mitigate bugs and achieve a repair success rate of 29.2% after two iterations, indicating substantial potential for LLMs to handle more complex problems. Our comprehensive and extensive study provides insights into the current limitations of LLM-based code generation and opportunities for enhancing the accuracy and quality of the generated code.
keywords:
@@ Abstract
1 Introduction
@@
Code generation (or program synthesis) aims to accurately generate executable programs from problem specifications such as natural language requirement descriptions and unit test examples [1, 2]. Research in code generation has gained considerable attention due to its significant impact on the software industry, particularly in enhancing productivity and accessibility by accelerating program development and improving code reliability [3, 4, 5].
@@
Recently, the advancement of large language models (LLMs) has significantly propelled the field of code generation [6, 7, 8, 9]. Trained on large corpora of text and code, LLMs can understand human requirements and generate the corresponding code through “next code token prediction” [10, 11, 12]. However, the existing LLMs still struggle to generate correct code when the requirements are complex [13, 14, 15, 16]. Despite the tremendous efforts made to enhance code generation capabilities via different means, such as using diverse pre-trained code corpora [17, 18, 19], fine-tuning models on higher-quality code [20, 21], or employing various natural language processing (NLP) techniques [22, 23, 24, 25], there remains a lack of a comprehensive understanding on the code generated by LLMs. Thus, an extensive large-scale empirical study is demanding, which can bring the following benefits.
First, comprehensive multi-dimensional code analysis beyond accuracy. Accuracy (i.e., passing rate of unit tests) is by far the most dominant code generation criterion, encompassing both syntactic and semantic requirements [26, 27]. Other characteristics, such as code conciseness and complexity, are equally significant due to the importance of code maintenance [28, 29]. Therefore, we expect the code generated by LLMs to be precise, concise, and low code complexity. Second, systematic bug taxonomy for LLM-generated code. Annotating the bug types in generated code and analyzing bug distributions are crucial for comprehensively understanding the limitations of LLMs in code generation. Specifically, previous research has primarily focused on superficial bugs in LLM-generated code, such as compilation failure and unit test failure. Yet, there is an absence of fine-grained categorization and investigation into the underlying causes of bugs in the generated code. Third, real-world benchmark construction and comparative analysis. There is a noticeable deficiency in benchmarks for evaluating LLMs in the real world. The benchmarks commonly used are constructed from programming questions [30, 15, 31] or derived from real-world code repositories and libraries [32, 33]. These repositories and libraries are also often used for pre-training LLMs, which could cause data contamination and test-case leakage [34, 35, 36, 31, 37, 38]. Therefore, developing a high-quality, previously untrained benchmark to evaluate LLMs under real-world conditions is crucial. Finally, practical bug mitigation method with empirical validation. While identifying bugs provides valuable insights, translating these findings into actionable solutions is essential. A method that enables LLMs to systematically critique and iteratively refine their own code, guided by comprehensive bug taxonomy and compiler feedback, could significantly enhance code generation quality without requiring additional model training.
@@
In this paper, we conduct an extensive large-scale empirical study to explore the capabilities and limitations of code generation based on LLMs. In particular, we thoroughly examine the correctness and complexity of the generated code by nine state-of-the-art LLMs. We then categorize bugs, and statistically model their distribution via in-depth analysis. To assess the effectiveness of LLMs in real-world code generation, we manually construct a high-quality benchmark, RWPB, and analyze the bugs in code generation based on it. Based on our findings, we propose a novel, practical method introducing self-critique to iteratively identify and fix bugs in LLM-generated code without additional training. Our study aims to answer the following research questions:
@@
  • RQ2: Bugs in Code Generated by LLMs (§4). What are the different types of root causes (i.e., bugs) of the LLM-generated code? How are these bugs distributed across widely used LLMs and popular benchmarks?
@@
  • RQ3: Benchmark Construction and Evaluation (§5). What is the effectiveness of LLMs in code generation on real-world projects? How do real-world project bugs differ from common, widely used benchmarks? Do the bug distributions generated on real-world projects differ from those on existing benchmarks?
@@
  • RQ4: Mitigating Bugs in Generated Code (§6). How to fix code bugs based on our findings without introducing additional resource consumption, such as fine-tuning LLMs.
@@
To investigate these RQs, we collected 1,163 programming problems from three widely used benchmarks (HumanEval+ [31], MBPP+ [30], and APPS+ [15]). To perform code generation, we selected nine LLMs, including three state-of-the-art closed-source LLMs (GPT-4 [39], GPT-3.5 [40], and Claude-3 [41]), four well-known open-source LLMs (DeepSeek-V3 [42], DeepSeek-R1 [43], Llama-3-Instruct [44] and Phi-3-Instruct [45]), and two popular open-source code LLMs (StarCoder-2 [18], DeepSeekCoder [17]). We evaluated these LLMs on generated code by using unit tests. Experimental results show that these LLMs could achieve on average a passing rate of 48.9%. Specifically, DeepSeek-R1 and DeepSeek-V3 demonstrate the best performance, which is 71.2% and 62.6%, respectively, while Phi-3 has the lowest accuracy at 30.9%.
@@
Furthermore, we collected the token number of problem descriptions; the length, the cyclomatic complexity and the API number of the canonical solution. This investigation allows us to analyze how the length of problem descriptions and the complexity of canonical solutions affect each model’s performance. We also examined the characteristics of the generated code, focusing on factors such as code characteristics and number of comments across different LLMs and benchmarks.
@@
Additionally, to explore the root causes of incorrect code, we conducted a comprehensive study on the bug taxonomy in such code. Specifically, our analysis includes two key steps: (1) First, we automatically created an initial taxonomy of potential bugs based on compiler feedback. (2) Second, human experts manually examined the incorrect code and annotated the bug type. During this phase, experts were given the flexibility to expand and refine the bug taxonomy as needed. In particular, this process involved 22 human experts with backgrounds in software engineering, with a minimum of two experts cross-verifying each annotation. Ultimately, we categorized the bugs into three primary and ten secondary types. Our study reveals several interesting findings: (i) Closed-source models surpass open-source models, particularly in handling complex problems. (ii) Functional bugs constitute the highest proportion, whereas syntax bugs represent the lowest. (iii) The code generated by LLMs typically has fewer lines but higher complexity and a similar API number compared to the canonical solutions. (iv) Although generating more comments during code generation can enhance LLMs’ reasoning abilities, LLMs still struggle to solve more complicated problems. (v) For complex problems, LLMs often fail to generate the optimal algorithms, thus leading to timeout errors.
@@
To evaluate the efficacy of LLM-based code generation in real-world projects, we developed a real-world benchmark by manually collecting 140 programming tasks from GitHub repositories established in 2024. Our experimental results revealed that closed-source LLMs outperform open-source LLMs in code generation quality for real-world projects, particularly excelling in reducing syntax bugs and runtime bugs. DeepSeek-R1 achieves the best performance with an accuracy of 77.9%, while Phi-3 has the lowest accuracy at only 22.1%. We also found that the bug distributions are inconsistent between the existing commonly used benchmarks and the real-world projects. Our study revealed that existing LLMs are prone to introducing bugs in code generation due to their failure to capture and interpret human intent and perform the correct corresponding logic, as evidenced in both current popular benchmarks and real-world projects. To mitigate these bugs, we propose an interactive approach involving self-critique [46] without additional training costs. Specifically, we make LLMs critique their incorrect code based on our constructed bug taxonomy and compiler feedback (i.e., error messages from failed executions) to understand the problem in-depth and analyze the causes of the bug, and then re-correct the code. This method can be performed iteratively. Experimental results showed that our proposed method can mitigate bugs and achieve a repair success rate of 29.2% after two iterations, which indicates substantial potential for LLMs to handle more complex tasks. Future improvements might focus on enhancing code quality from an artificial intelligence perspective (e.g., improving LLM comprehension and mitigating the misuse of function APIs by introducing explanatory comments to training code) and from a software engineering standpoint (e.g., preventing syntax bugs by introducing additional syntax checkers rather than solely relying on a conventional generation paradigm in the inference time).
@@
  • A thorough evaluation of the capabilities of LLMs in code generation. We conduct extensive experiments to evaluate the performance of current widely used LLMs. We also explore the effect of different types of tasks on code generation and the characteristics of generated code.
  • A bug taxonomy of LLM-generated code. We develop a bug taxonomy based on code generated by LLMs. We further analyze these bugs and model their distribution to understand the root causes of inaccuracies.
  • A real-world benchmark. We manually construct a real-world benchmark, RWPB, by collecting code from the newest GitHub repositories. We also analyze and compare the distribution of bugs between existing benchmarks and our real-world benchmark, highlighting the inadequacies of current models in real-world scenarios.
  • A novel method to mitigate code generation bugs. We propose a novel approach to mitigate bugs by introducing self-critique: allowing LLMs to find and fix bugs by iteratively critiquing their generated code. We also offer some suggestions and possible future solutions for improvement.
2 Experimental Design
Large Language Model Selection. To thoroughly evaluate LLM-based code generation, we selected nine large language models of three types: open-source text LLMs, open-source code LLMs, and closed-source LLMs. For open-source text LLMs, we selected four common models, namely DeepSeek-V3 [42], DeepSeek-R1 [43], Llama-3 [44] and Phi-3 [45]. DeepSeek-V3 is a Mixture-of-Experts language model that achieves high performance while being significantly more cost-efficient to train and deploy. DeepSeek-R1 is a reasoning-specialized language model that is trained through reinforcement learning and chain-of-thought training based on DeepSeek-V3. Llama-3 is an advanced and popular LLM developed by Meta-AI, which is trained on extensive text and code corpora. Phi-3 is another extensively used LLM, noted for its training on high-quality textbook datasets. In the category of open-source code LLMs, our study used StarCoder-2 [18] and DeepSeekCoder [17]. StarCoder-2 is a widely used code model trained on 3.3 to 4.3 trillion tokens of code from The Stack v2 dataset [18], encompassing over 600 programming languages. DeepSeekCoder is another state-of-the-art open-source code LLM trained on 2 trillion code and natural language tokens. We also selected three universally used closed-source LLMs, namely GPT-3.5 [40], GPT-4 [39], and Claude-3 [41]. We set the temperature to 0.1 and use top-1 sampling to ensure consistent results. Additional information and details on these models are provided in Table 1.
Table 1: Overview of nine LLMs used in our study.
Modality Model Context Window Release Date
@@ 2 Experimental Design
Benchmark Selection. To evaluate the performance of selected LLMs and analyze bugs in code generation, we constructed a collection of assessments from well-known benchmarks. Our selection criteria included widely recognized benchmarks to ensure relevance, the exclusion of popular libraries or projects to avoid data leakage, and diverse unit tests with high code coverage. The target programming language for the generated code was set to Python, as most current research focuses on improving code generation in Python [17, 47, 48]. Finally, we selected three widely used benchmarks, HumanEval+ [31], MBPP+ [30] and APPS+ [15]. In particular, HumanEval is a popular interview Python benchmark, designed to test language comprehension, algorithmic thinking, and basic mathematics skills. MBPP is another prevalent introductory Python benchmark created by crowd-sourced developers. Both HumanEval+ and MBPP+ are expansions of their original versions, augmenting additional unit tests and comprising 164 and 399 programming problems, respectively [49]. APPS is a Python benchmark, including three difficulty levels (introductory, interview, and competition), collected from open-access sites, including Codewars [50], AtCoder [51], Kattis [52], and Codeforces [53]. APPS+ is a curated version of APPS that includes manual analysis to eliminate issues such as incomplete code and syntax errors, API misuse, and missing library dependencies [13]. For APPS+, we selected 200 samples each from the categories to maintain difficulty diversity. Our benchmark includes 164 samples from HumanEval+, 399 samples from MBPP+, and 600 samples from APPS+, totaling 1,163 samples. The average code coverage is 96.0% as shown in Table 2.
@@
Additionally, to evaluate the efficacy of LLM in real-world projects, we created a real-world benchmark from GitHub repositories established in 2024 (details in Section 5.1).
@@
Prompt Design. To ensure experimental consistency across various LLMs, we used a standardized prompt format based on prior studies [49, 13]. Notably, the prompt excluded library import information, requiring LLMs to autonomously identify and import necessary libraries for the APIs used. The prompt can be found in our artifact repository [54].
@@
Computation and Labor Resources. To evaluate current LLMs for code generation and validate our proposed improvements, we conducted experiments on a single node with eight A100-80GB GPUs and 2TB of CPU memory, using Python 3.11. For code evaluation, bug analysis, and the construction of our real-world project benchmark RWPB, we assigned 37 annotators in software engineering. Given their countries of residence, all annotators obtained adequate compensation.
3 LLM-Based Code Generation